OpenStudy (anonymous):

what is the ans for {(-1)^2+(-2)^3+(-3)^4} divided by {(-1)^4+(-2)^3+(-3)^2 ??

OpenStudy (mathteacher1729):

Question: Do you know how to take powers of negative numbers?

OpenStudy (anonymous):

yeah

OpenStudy (mathteacher1729):

Ok. So let's take this apart piece by piece. First the numerator (top of the fraction) What is (-1)^2 = (-2)^3 = (-3)^4 =

OpenStudy (anonymous):

(-1)^2 is 1 over 1, (-2)^3is 2 over 8 and (-3)^4 is 3 over 81

OpenStudy (mathteacher1729):

There are no fractions yet. (-1)^2 = 1 (-2)^3 = -8 (-3)^4 = 81

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

ooooo i got it! thankx a lot...... :)

OpenStudy (mathteacher1729):

Now do that for the denominator. When you're done, add (or subtract) the numbers in the numerator, then add (or subtract) the numbers in the denominator. Reduce the fraction, and you're done. :)

OpenStudy (anonymous):

thankxxxxxxxxxxxxxxxxxx